OpenStudy (anonymous):

After the softball game, Alise's coach bought each of the 14 players a double-scoop or triple-scoop ice cream cone at the local ice cream shop. The coach bought a total of 37 scoops of ice cream. Use a model to determine how many girls got a triple-scoop ice cream cone. A. 6 B. 7 C. 8 D. 9

OpenStudy (anonymous):

6

OpenStudy (anonymous):

This can be modeled by simultaneous equations. Let x = number of girls with double scoops, y = number of girls with triple scoops. x+y=15 (because 15 girls total) x = 15-y 36 = 2x+3y (36 scoops, 2 scoops per double girl, 3 scoops per triple girl) 36 = 2(15-y)+3y (by substituting 15-y for x) 36 = 30-2y+3y 36 = 30+y y = 6

OpenStudy (anonymous):

Thank you :-)

OpenStudy (anonymous):

no thank you for all your help eailer

OpenStudy (anonymous):

Lol your welcome :-)